### v1

# Demo: Soccer Information Retrieval via Natural Queries using SoccerRAG

Abstract
The rapid evolution of digital sports media necessitates sophisticated information retrieval systems that can efficiently parse extensive multimodal datasets. This paper demonstrates SoccerRAG, an innovative framework designed to harness the power of Retrieval Augmented Generation (RAG) and Large Language Models (LLMs) to extract soccer-related information through natural language queries. By leveraging a multimodal dataset, SoccerRAG supports dynamic querying and automatic data validation, enhancing user interaction and accessibility to sports archives. We present a novel interactive user interface (UI) based on the Chainlit framework which wraps around the core functionality, and enable users to interact with the SoccerRAG framework in a chatbot-like visual manner.
I Introduction
The burgeoning interest in Retrieval Augmented Generation (RAG), fueled by the rapid advancements in Large Language Models (LLMs), has paved new avenues for exploring innovative use cases across various domains, notably within multimodal information retrieval [1]. Despite the wide applicability of RAG frameworks in enhancing the capabilities of generative AI for open-domain question answering and beyond, its potential in the sports domain, particularly within soccer analytics, remains largely untapped. Our recent research has endeavored to bridge this gap by harnessing multimodal datasets inherent to the soccer broadcast pipeline and integrating RAG with different data modalities. We have presented SoccerRAG, a framework for retrieving multimodal soccer information using natural language queries, from an augmented soccer dataset based on SoccerNet [2, 3], which includes game videos with image frames and audio, timestamped captions (transcribed audio), annotations for game events, and player information.
In this paper, we demonstrate the use of the SoccerRAG framework through the command-line interface (CLI) and a novel user interface (UI). Figure 1 presents an overview of the SoccerRAG framework, with the core components (database, feature extractor, feature validator, and SQL agent) as presented earlier, wrapped by a novel interactive UI which serves as a user-friendly tool for visual interactions. The UI was created using the Chainlit framework. Chainlit is an open-source Python package that enables developers to build production-ready conversational AI applications, providing features for quick integration. It offers integrations with popular libraries and frameworks, such as OpenAI and LangChain, and allows for custom front-ends with React-based user interfaces [4]. The SoccerRAG integrates functionality for data representation, feature extraction and validation, database querying (based on our proposed database schema for the augmented version of the SoccerNet dataset), and a novel extractor-validator chain. The open source implementation for SoccerRAG is accessible under [5] and fully reproducible through the instructions provided therein, which are detailed below.
Figure 1: Overview of the SoccerRAG framework, including the novel interactive UI.
The potential of our proposed system for information retrieval offers a promising avenue for fans and broadcasters alike, enabling the recollection of game specific information and highlights, as well as broader statistics and insights (within the scope of the entire database) through natural queries. Pursuing this avenue of research, we aspire not only to contribute to the body of knowledge surrounding RAG and multimodal information retrieval, but also to pioneer a smart system that encapsulates the dynamic essence of soccer, fostering a deeper connection between the sport and its global audience.
II Demonstration
II-A Application Flow
The core functionality of the SoccerRAG framework can be represented with the following application flow:
  • The user provides a natural language query related to the contents of the database.
  • The user input is sent to an LLM along with the properties schema and a system prompt describing the properties the LLM should extract from the query. The LLM then returns a list of extracted properties relevant to the query.
  • Each extracted feature is checked against the appropriate table in the database using string matching algorithms. This step aims to correct spelling mistakes and abbreviations. Once a value is found, both the value and its primary key are added to the extracted value.
  • The cleaned user prompt is combined with system-specific prompts to guide the LLM in generating SQL queries that will answer the user’s natural language query. The constructed query is then passed to the SQL chain, which designs and executes the SQL queries against the underlying database. The SQL chain handles communication between the system and the database, retrieves the requested data, and prepares the results for presentation to the user.
II-B Artifacts
Figure 2 presents the structure of the SoccerRAG codebase, publicly accessible under [5]. The data folder is used by the core pipeline to read in the source dataset (including the original SoccerNet data, as well as our augmented league and team tables), and storing the database files. The src folder contains the configuration and executable files. The root directory includes the main executable (main.py), environment variables (.env), and a list of dependencies (requirements.txt). A proof-of-concept deployment of the SoccerRAG UI can be found under: https://simulamet-host-soccerrag.hf.space/
G SoccerRAG
? data/
{ dataset/
0 augmented_leagues.csv
0 augmented_teams.csv
<downloaded assets>
/ src/
/ database.py
/ extractor.py
/ sql_chain.py
2 conf/
schema.json
sqls.json
.env
/ app.py
/ main.py
/ main_cli.py
requirements.txt
/ setup.py
Figure 2: Structure of the SoccerRAG codebase.
II-C Requirements
The SoccerRAG codebase requires Python 3.12 or above to be run. The requirements.txt file in the codebase specifies the required packages, which can be installed through pip. SoccerRAG does not require a GPU, and can be run on any machine with a CPU. For the experiments presented in the research paper, a machine with a Windows 11 operating system, Intel(R) Core(TM) i5-9300H CPU @2.40GHz, and 16GB memory has been used. SoccerRAG requires an OpenAI API key, which should be specified in the .env file. In order to log all queries to the OpenAI endpoints, it is possible to set a LangSmith API key [6]. This makes it possible to monitor the cost of each call, and keep history.
II-D Running the Framework
Below are the steps to run the SoccerRAG framework (core functionality described in Section II-A).
II-D1 Configuration
The SoccerRAG framework is set up using environment variables. Table I presents an overview of the configuration parameters. The parameters marked as mandatory do not have default values, and need to be configured before runtime. There should be an entry in .env for each configuration parameter. The SoccerRAG framework has been tested with GPT-3.5-Turbo [7] and GPT-4.0-Turbo [8, 9], but can support any model available from OpenAI [10]. The model version can be specified through the OPENAI_MODEL configuration parameter.
Parameter Description M Default
OPENAI_API_KEY OpenAI API key NA
OPENAI_MODEL GPT version gpt-3.5-turbo-0125
DATABASE_URL Location of DB "/data/games."
LANGSMITH Trace api calls False
LANGSMITH_API_KEY API key for LangSmith  [^1] If LANGSMITH is set to True. NA
LANGSMITH_PROJECT Project tag for LangSmith SoccerRag
FEW_SHOT Num. options in user validation 3
TABLE I: SoccerRAG configuration parameters (M: mandatory).
II-D2 Running via UI
Run python setup.py to download the dataset and set up the DB, and chainlit run app.py for the UI.
(a) Black background
(b) White background
Figure 3: SoccerRAG UI example 1. The user prompts: "Can you calculate Real Madrids [sic] home advantage for the 2015/16 season?"
Figure 3 presents an example screenshot, depicting alternative background colors [^2] SoccerRAG UI background color is configurable: click the user icon in the top right corner, select settings, then toggle the dark mode switch. for the UI, as well as the intermediate feedback to users that the SoccerRAG provides for each operational step, which are retrieved from the feature extractor and feature validator components in the core pipeline. The meaningful response returned to the user prompt "Can you calculate Real Madrids [sic] home advantage for the 2015/16 season?" demonstrates that the framework is able to grasp soccer-related concepts such as a home advantage (which refers to the difference in scores achieved by a team in home vs. away games).
Figure 4: SoccerRAG UI example 2. The user prompts: "How many yellow cards did messi [sic] get in the 2015-16 season on home turf?"
Figures 4 and 5 present further example screenshots, demonstrating the capabilities of SoccerRAG which range from database-wide statistics (example 2: "How many yellow cards did messi [sic] get in the 2015-16 season on home turf?" is returned with the correct aggregation across all games of the desired season), to supporting multi-format information presentation (example 3: "List all games played by ManU [sic] in the 16-17 season. / Give the result as a markdown table with following format / HomeTeam AwayTeam Score Venue Attendance Date").
Figure 5: SoccerRAG UI example 3. The user prompts: "List all games played by ManU [sic] in the 16-17 season. / Give the result as a markdown table with following format / HomeTeam AwayTeam Score Venue Attendance Date"
One of the challenges faced by systems which are prompted via natural language queries is the potential lack of clarification with respect to the entities in the prompt. In the context of soccer, these can be player, team, or league names, places, season dates, etc. SoccerRAG requests clarification from the user in cases of inadequate information, as shown in Figure 6. Here (example 4: "Create a list of all games Lionel got a yellow card / Make the list in markdown with following coulms [sic] / GameId, HomeTeam, AwayTeam, Score, Date"), as opposed to directly returning a response ("one-step") as in examples 1-3, the framework needs to ask the user to select the correct value for "Lionel", as there are multiple entities in the database with this name ("two-step"). In the UI, the user can click a button to communicate their choice and validate one of the presented options, or pass the original string untranslated to the next component.
(a) Step 1: user asked for clarification
(b) Step 2: response returned based on additional user input
Figure 6: SoccerRAG UI example 4. The user prompts: "Create a list of all games Lionel got a yellow card / Make the list in markdown with following coulms [sic] / GameId, HomeTeam, AwayTeam, Score, Date".
II-D3 Running via CLI
[…]
Figure 7: SoccerRAG CLI example 1. The user prompts: "In the game between Chelsea and Burnley in the 2014-15 season, did anyone get a yellow card? If yes, who."
Figure 7 presents an example screenshot, depicting the console outputs from the CLI for the query "In the game between Chelsea and Burnley in the 2014-15 season, did anyone get a yellow card? If yes, who.", which receives a response directly (one-step).
[…]
III Conclusion
We demonstrate SoccerRAG, a framework that leverages RAG and LLMs to efficiently retrieve multimodal soccer information via natural language queries, enhancing the accessibility of sports datasets by allowing for intuitive user interactions with complex data archives. In addition to the core pipeline presented in our research paper, we have introduced a novel UI to facilitate visual user interactions with the framework in a chatbot-like manner. The UI serves as a wrapper around the core pipeline, whereby it passes the user query to the feature extractor component, after which the query is validated by the feature validator against the database, the query with validated properties are passed to SQL agent, which retrieves data from the database and returns a response, and the UI presents the response to the user.
We present the entire codebase for the SoccerRAG as open source software, along with instructions to run the framework either via CLI or via UI. Our experiments up to now demonstrate SoccerRAG’s capability to accurately interpret complex queries and facilitate dynamic user engagements with soccer content. We believe that SoccerRAG exemplifies the transformative potential of AI in sports analytics, promising enriched global fan experiences and advances in the accessibility of sports information.
Acknowledgment
This research was partly funded by the Research Council of Norway, project number 346671 (AI-storyteller).